# Variational Learning of Disentangled Representations

### v1

@@ E.1 Considerations and reproducibility
Below, we provide additional details for the hyperparameters used in each experiment, and any other external resources used to obtain the corresponding sections’ results.
E.1.1 Parametric Model
@@
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
@@
$w$ Stimulation ($\uparrow$) $z$ Cell Type ($\uparrow$) $z$ Stimulation ($\downarrow$)
@@
CSVAE $0.939\pm 0.002$ $0.406\pm 0.001$ $\mathbf{0.002\pm 0.0}$
HCSVAE - No Adv. $0.933\pm 0.006$ $0.628\pm 0.016$ $0.091\pm 0.002$
HCSVAE $0.931\pm 0.005$ $0.433\pm 0.001$ $0.003\pm 0.0$
DIVA $0.801\pm 0.0$ $0.628\pm 0.011$ $0.056\pm 0.0$
@@
DISCoVeR (Ours) $\mathbf{0.946\pm 0.002}$ $\mathbf{0.688\pm 0.0}$ $\mathbf{0.002\pm 0.0}$
E.1.2 Noisy Swiss Roll
@@
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
@@
DISCoVeR (Ours) 0.9 0.2 0.2 8 0.1
E.1.3 Noisy Colored MNIST
For this experiment, we consider $z\in\mathbb{R}^{20}$, $w\in\mathbb{R}^{2}$ and use convolutional neural networks (CNNs) to parameterize approximate posteriors and the generative model. For this example, DISCoVeR can support $z,w$ with different sizes, by parameterizing $p(w\mid y)$ through neural networks. For all models, we use learning rate $\gamma=0.0001$. We detail the architectures and model-specific hyperparameters in Supplementary Tables 7 - 10. All other neural networks are formulated as MLPs with $n_{hidden}=2,d_{hidden}=4096$.
@@ 3 Linear(2)
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
@@
DISCoVeR (Ours) 0.5 0.0001 0.0001 0.1 0.5
E.1.4 CelebA-Glasses
Motivated by the previous application of Klys et al. [2018], our choices follow those outlined in Larsen et al. [2016]. We provide a detailed table of model-specific hyperparameters in Supplementary Table 11:
@@ E.1.4 CelebA-Glasses
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
@@
DISCoVeR (Ours) 1000000 0.0001 0.0001 2000 100000
E.1.5 scRNA-Seq
Following on the previous applications by Lopez et al. [2018], we use $z\in\mathbb{R}^{10}$, $w\in\mathbb{R}^{2}$. For DISCoVeR  we match the dimensions and use $w\in\mathbb{R}^{10}$ to avoid parameterizing the prior $p(w\mid z,y)$ with an additional neural network. We use MLPs with $n_{hidden}=1,d_{hidden}=128$ to parameterize approximate posteriors, the generative model and classifiers. We calculate K-Means NMI through scikit-learn [Pedregosa et al., 2011] by calling the normalized_mutual_info_score function with the original labels and the clusterings obtained by running KMeans on (1) the entire latent embedding and (2) single dimensions of the embedding and report the highest score. A more detailed table of model-specific hyperparameters is provided in Supplementary Table 12:
@@
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
@@
DISCoVeR (Ours) 0.9 0.0001 0.0001 100 0.1
E.2 Summary of the scVI generative model for 5.2.3
@@
Here, $g$ indexes genes, $l_{i}=\sum_{g}x_{ig}$ denotes the total number of counts for a single cell, $z_{i}$ denotes the latent representation of the cell, and $\rho_{i}$ denotes the normalized expression of the cell. $f_{\theta}$ is formulated as a neural network with a final softmax layer. $h_{\phi}$ is a neural network used to parameterize zero-inflation probabilities for the generative zero-inflated negative binomial (ZINB) distribution. As such, for a single batch, the formulation of scVI is equivalent to the VAE with a ZINB likelihood. While all other models can be extended easily, DISCoVeR requires reconstructions as a proxy for the adversarial loss. For this formulation, we directly treat the normalized expressions $\rho_{i}$ as the adversarial reconstructions $\hat{x}$.

### v2

@@ H.1 Considerations and reproducibility
Below, we provide additional details for the hyperparameters used in each experiment, and any other external resources used to obtain the corresponding sections’ results. In addition, we include details regarding runtime and memory footprint of running experiments with the models included in our study.
Supplementary Table 13: Time spent per epoch during training for each dataset.
P.M. N.S.R CMNIST CelebA scRNA-seq
CSVAE - N.A. 10.91s 8.02s 14s 54.43s 4.78s
CSVAE 12.71s 8.98s 14.41s 74.68s 4.99s
HCSVAE - N.A. 15.6s 10.92s 17s 44.3s 6.34s
HCSVAE 16.51s 11.8s 16.8s 68.53s 5.3s
DIVA 11.1s 8s 18.59s 48.96s 3.55s
CCVAE 12.1s 8.44s 17.9s 49.65s 5.25s
DisCoVR(Ours) 12.18s 8.73s 21.8s 109.86s 6.09s
Supplementary Table 14: Model inference time for a single batch for each dataset.
P.M. N.S.R CMNIST CelebA scRNA-seq
CSVAE - N.A. 72ms 28ms 57ms 329ms 100ms
CSVAE 40ms 29ms 59ms 187ms 96ms
HCSVAE - N.A. 36ms 30ms 55ms 214ms 95ms
HCSVAE 37ms 39ms 52ms 191ms 119ms
DIVA 25ms 26ms 60ms 154ms 42ms
CCVAE 27ms 28ms 42ms 163ms 93ms
DisCoVR(Ours) 32ms 27ms 63ms 205ms 123ms
Supplementary Table 15: Memory footprint of running an experiment for each dataset.
P.M. N.S.R CMNIST CelebA scRNA-seq
CSVAE - N.A. 53 MiB 255 MiB 1988 MiB 4868 MiB 298 MiB
CSVAE 253 MiB 255 MiB 2378 MiB 4812 MiB 300 MiB
HCSVAE - N.A. 254 MiB 255 MiB 2558 MiB 4588 MiB 292 MiB
HCSVAE 253 MiB 256 MiB 2998 MiB 4466 MiB 294 MiB
DIVA 253 MiB 255 MiB 2634 MiB 4996 MiB 300 MiB
CCVAE 253 MiB 255 MiB 3066 MiB 4998 MiB 300MiB
DisCoVR (Ours) 254 MiB 257 MiB 3612 MiB 7078 MiB 308MiB
H.1.1 Parametric model
@@
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
@@
$w$ - Stimulation ($\uparrow$) $z$ - Cell Type ($\uparrow$) $z$ - Stimulation ($\downarrow$)
@@
[…]
HCSVAE - No Adv. $0.933\pm 0.006$ $628\pm 0.016 0.0$$091\pm 0.002$
HCSVAE $0.931\pm 0.005$ $433\pm 0.001 0.0$$003\pm 0.0$
DIVA $0.801\pm 0.0$ $628\pm 0.011 0.0$$056\pm 0.0$
@@
DisCoVR (Ours) $\mathbf{0.946\pm 0.002}$ $\mathbf{0.688\pm 0.0}$ $\mathbf{0.002\pm 0.0}$
H.1.2 Noisy Swiss roll
@@
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
@@
DisCoVR (Ours) 0.9 0.2 0.2 8 0.1
H.1.3 Noisy colored MNIST
For this experiment, we consider $z\in\mathbb{R}^{20}$, $w\in\mathbb{R}^{2}$ and use convolutional neural networks (CNNs) to parameterize approximate posteriors and the generative model. For this example, DisCoVR can support $z,w$ with different sizes, by parameterizing $p(w\mid y)$ through neural networks. For all models, we use learning rate $\gamma=0.0001$. We detail the architectures and model-specific hyperparameters in Supplementary Tables 19 - 22. All other neural networks are formulated as MLPs with $n_{hidden}=2,d_{hidden}=4096$.
@@ 3 Linear(2)
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
@@
DisCoVR (Ours) 0.5 0.0001 0.0001 0.1 0.5
H.1.4 CelebA-glasses
Motivated by the previous application of CSVAE, our choices follow those outlined in pmlr-v48-larsen16. We provide a detailed table of model-specific hyperparameters in Supplementary Table 23:
@@ H.1.4 CelebA-glasses
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
@@
DisCoVR (Ours) 1000000 0.0001 0.0001 2000 100000
H.1.5 scRNA-Seq
Following on the previous applications by scVI, we use $z\in\mathbb{R}^{10}$, $w\in\mathbb{R}^{2}$. For DisCoVR  we match the dimensions and use $w\in\mathbb{R}^{10}$ to avoid parameterizing the prior $p(w\mid z,y)$ with an additional neural network. We use MLPs with $n_{hidden}=1,d_{hidden}=128$ to parameterize approximate posteriors, the generative model and classifiers. We calculate K-Means NMI through scikit-learn (scikit-learn) by calling the normalized_mutual_info_score function with the original labels and the clusterings obtained by running KMeans on (1) the entire latent embedding and (2) single dimensions of the embedding and report the highest score. A more detailed table of model-specific hyperparameters is provided in Supplementary Table 24:
@@
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
@@
DisCoVR (Ours) 0.9 0.0001 0.0001 100 0.1
H.2 Summary of the scVI generative model for 5.2.3
@@
Here, $g$ indexes genes, $l_{i}=\sum_{g}x_{ig}$ denotes the total number of counts for a single cell, $z_{i}$ denotes the latent representation of the cell, and $\rho_{i}$ denotes the normalized expression of the cell. $f_{\theta}$ is formulated as a neural network with a final softmax layer. $h_{\phi}$ is a neural network used to parameterize zero-inflation probabilities for the generative zero-inflated negative binomial (ZINB) distribution. As such, for a single batch, the formulation of scVI is equivalent to the VAE with a ZINB likelihood. While all other models can be extended easily, DisCoVR requires reconstructions as a proxy for the adversarial loss. For this formulation, we directly treat the normalized expressions $\rho_{i}$ as the adversarial reconstructions $\hat{x}$.